OpenStudy (anonymous):

Solve this equation. -2 log8(x) = log8(49) and 8 are the bases

OpenStudy (anonymous):

log8(x^-2)=log8(49) -2*log8(x)=2*log8(7) -log8(x)=log8(7) log8(1/x)=log8(7) 1/x=7 x=1/7